chuckthedonut:

On a map, ½ inch represents 4/5 mile. What distance on the map represents 1 mile? _____ inches on map = 1 mile Show your work.

chuckthedonut:

1 attachment
jhonyy9:

so you can use the rule of 3 simple and in this way you write \[\frac{ 1 }{ 2 } inch = \frac{ 4 }{ 5 } mile\] x inches = 1 mile cross multiplie and get \[\frac{ 1 }{ 2 } = \frac{ 4 }{ 5 } x\] cross multiplie and get 5 = 8x divide by 8 \[\frac{ 5 }{ 8 } = x\] so in this way result 5/8 inches on map = 1 mile - the proof : 1/2 = 4/5 5/8 = 1 1/2 *1 = 4/5 *5/8 1/2 = 20/40 1/2 = 1/2 hope you understand my above calcules